### v2

@@ 1.1 Gameability Study: Finding Extreme Maps that aren’t Out of Bounds
We define a map as “extreme” using the most straightforward of terms: it results an extreme number of districts won by a given party, while still maintaining the basic redistricting criteria of population balance, contiguity and some measure of compactness. Thus, in order to determine whether a metric is “gameable,” we must determine if extreme maps (maps with a high number of districts won by a particular party) are achievable while keeping the metric within acceptable bounds. In the case of the Efficiency Gap, the creators were careful not to prescribe the “acceptable bounds”. However, they used a bound of $\pm 8\%$ for the bound between “presumptive validity and invalidity” in their paper introducing the metric [PartisanGerrymanderingEfficiencyGap]. For consistency, we use the corresponding values here for all of our metrics: $\pm 16\%$ for metrics whose range is $\left[-1,1\right]$ (the GEO Metric Ratio [^2] A value obtained using the two GEO metric scores, which we define later. and Declination), and $\pm 8\%$ for metrics whose range is $\left[-0.5,0.5\right]$ (the Efficiency Gap and Mean-Median Difference). While we needed to fix a consistent range of metric values in order to complete our Gameability Study, in Section LABEL:sec:conclusions we explain why this choice does not impact our ultimate conclusions.
@@ 1.2 Extreme Metric Values versus Extreme Maps
In Section LABEL:subsec:results_no_restriction, we discuss the range of metric values on maps with the maximum or near-maximum number of Democratic-won (or Republican-won) districts. We consider the same metrics that we considered for the Gameability Study of Section 1.1: the Mean-Median Difference, Efficiency Gap, Declination, and GEO metric. We shall see that extreme values of the Mean-Median Difference do not correspond to maps with an extreme number of districts won (although extreme values of the other metrics do correspond to maps with an extreme number of districts won).
@@ 1.3 Metrics and Ensembles
Using ensemble analysis on “number of districts won” by a particular party is a straightforward comparison when it comes to gerrymandering: we are determining whether the proposed map has an extreme number of districts won, compared to the other maps in the ensemble. The implications are easy to interpret. However, if we use ensemble analysis on another metric (such as the Mean-Median Difference, Efficiency Gap, Declination, and GEO metric) it is less clear how to interpret the results. Does an outlier Efficiency Gap value imply the map has an extreme number of districts won by some party? What about an outlier Mean-Median Difference value? Our results of this study are in Section LABEL:subsec:ensembles. In short, we see that the answer is “no” for the Mean-Median Difference (that is, outlier values of the Mean-Median Difference do not correspond to maps with an extreme number of districts won) and “yes” for the other metrics.
@@ 2 Metrics and Detecting Gerrymandering
As stated in Section 1, we will study four metrics which do not require detailed map data and are commonly used to detect partisan gerrymandering: the Mean-Median Difference, Efficiency Gap, Declination, and Geography and Election Outcome (GEO) metric. Each of these metrics assumes that there are two political parties vying for seats. We study these metrics both in terms of their gameability (results in Section LABEL:subsec:results_metrics_restricted) as well as whether extreme metric values correspond to extreme maps (results in Section LABEL:subsec:results_no_restriction). We also study how each of these metrics does or does not detect gerrymandering when evaluated on an ensemble of redistricting maps (results in Section LABEL:subsec:ensembles). For the reader unfamiliar with these metrics, we define them here. But first we highlight why we ground our understanding of the existance of gerrymandering in “districts won.”
@@
In our Gameability Study, we seek to determine whether a metric is “gameable” by determining if a metric can be within reasonable bounds, but allow a map with an extreme number of districts won by some party. In our Extreme Metric Values versus Extreme Maps Study, we look to see if extreme values of each metric correlate with “extreme maps” as defined in Section 1.1: an extreme number of districts won by some party. And in our study of Metrics and Ensembles, we consider a map to be likely gerrymandered if that map’s number of districts won is an outlier among maps in the ensembles (that is, if the party has an extreme number of districts won).
Thus, for each of these studies, we are grounding our understanding of gerrymandering in an extreme number of districts won (or predicted to be won by partisan data). This certainly follows the common conception of gerrymandering. For example, the Brennan Center for Justice, in their “Gerrymandering Explained” report [BrennanCenterGerrymanderingExplained], describe gerrymanders in terms of the number of additional seats a party won due to gerrymandering. Intuition and common understanding aside, Political Scientists and Social Scientists also describe gerrymandering in terms of seats won: “Partisan gerrymanderers use their knowledge of voter preferences and their ability to draw favorable redistricting plans to maximize their party’s seat share [626161].” Mathematicians and Computer Scientists agree with this definition: “[T]he express intent of a partisan gerrymander is to secure for their party as many seats as possible under the constraints of voter geography and the other rules of redistricting” [ImplementingPartisanSymmetry]. While there may be other factors that appear in a gerrymandered map, such as durability of the map over time or incumbency protection, we hold that the baseline goal of a gerrymander is to push the number of districts won towards extreme values for the favored party.
@@ 2.4 Declination
Let $F$ be the center of mass of points in $\mathcal{A}$, $G=\left(\frac{k}{n},\frac{1}{2}\right)$, and $H$ the center of mass of points in $\mathcal{B}$. We define angle $\theta_{P}$ to be the angle that the line through $F$ and $G$ makes with the line $y=\frac{1}{2}$, and $\theta_{Q}$ to be the angle that the line through $G$ and $H$ makes with the line $y=\frac{1}{2}$. These can be visualized in Figure 1, which is from from Warrington’s article [WarringtonDeclinationELJ]:

### v3

@@ 1.1 Gameability Study: Finding Extreme Maps that aren’t Out of Bounds
We define a map as “extreme” using the most straightforward of terms: it results in an extreme number of districts won by a given party, while still maintaining the basic redistricting criteria of population balance, contiguity and some measure of compactness. Thus, in order to determine whether a metric is “gameable,” we must determine if extreme maps (maps with a high number of districts won by a particular party) are achievable while keeping the metric within acceptable bounds. In the case of the Efficiency Gap, the creators were careful not to prescribe the “acceptable bounds”. However, they used a value of $\pm 8\%$ for the bound between “presumptive validity and invalidity” in their paper introducing the metric [PartisanGerrymanderingEfficiencyGap]. For consistency, we use the corresponding values here for all of our metrics: $\pm 16\%$ for metrics whose range is $\left[-1,1\right]$ (the GEO Metric Ratio [^2] A value obtained using the two GEO metric scores, which we define later. and Declination), and $\pm 8\%$ for metrics whose range is $\left[-0.5,0.5\right]$ (the Efficiency Gap and Mean-Median Difference). While we needed to fix a consistent range of metric values in order to complete our Gameability Study, in Section LABEL:sec:conclusions we explain why this choice does not impact our ultimate conclusions.
@@ 1.2 Extreme Metric Values versus Extreme Maps
In Section LABEL:subsec:results_no_restriction, we discuss the range of metric values on maps with the maximum or near-maximum number of Democratic-won (or Republican-won) districts. We consider the same metrics that we considered for the Gameability Study of Section 1.1. We shall see that extreme values of the Mean-Median Difference do not correspond to maps with an extreme number of districts won (although extreme values of the other metrics do correspond to maps with an extreme number of districts won).
@@ 1.3 Metrics and Ensembles
Using ensemble analysis on “number of districts won” by a particular party is a straightforward comparison when it comes to gerrymandering: we are determining whether the proposed map has an extreme number of districts won, compared to the other maps in the ensemble. The implications are easy to interpret. However, if we use ensemble analysis on another metric, it is less clear how to interpret the results. Does an outlier Efficiency Gap value imply the map has an extreme number of districts won by some party? What about an outlier Mean-Median Difference value? Our results of this study are in Section LABEL:subsec:ensembles. In short, we see that the answer is “no” for the Mean-Median Difference (that is, outlier values of the Mean-Median Difference do not correspond to maps with an extreme number of districts won) and “yes” for the other three metrics we consider.
@@ 2 Metrics and Detecting Gerrymandering
Each of the four metrics we study assumes that there are two political parties vying for seats. For completeness, we define these four metrics here, but first we highlight why we ground our understanding of the existence of gerrymandering in “districts won.”
@@
Gerrymandering is commonly understood as the manipulation of district boundaries to produce an extreme number of districts won (or predicted to be won) by a particular party. For example, the Brennan Center for Justice, in their “Gerrymandering Explained” report [BrennanCenterGerrymanderingExplained], describe gerrymanders in terms of the number of additional seats a party won due to gerrymandering. Intuition and common understanding aside, Political Scientists and Social Scientists also describe gerrymandering in terms of seats won: “Partisan gerrymanderers use their knowledge of voter preferences and their ability to draw favorable redistricting plans to maximize their party’s seat share [626161].” Mathematicians and Computer Scientists agree with this definition: “[T]he express intent of a partisan gerrymander is to secure for their party as many seats as possible under the constraints of voter geography and the other rules of redistricting” [ImplementingPartisanSymmetry]. While there may be other factors that appear in a gerrymandered map, such as durability of the map over time or incumbency protection, we hold that the baseline goal of a gerrymander is to push the number of districts won towards extreme values for the favored party.
@@ 2.4 Declination
Let $F$ be the center of mass of points in $\mathcal{A}$, $G=\left(\frac{k}{n},\frac{1}{2}\right)$, and $H$ the center of mass of points in $\mathcal{B}$. We define angle $\theta_{P}$ to be the angle that the line through $F$ and $G$ makes with the line $y=\frac{1}{2}$, and $\theta_{Q}$ to be the angle that the line through $G$ and $H$ makes with the line $y=\frac{1}{2}$. These can be visualized in Figure 1, which is from Warrington’s article [WarringtonDeclinationELJ]: